Status Report

ISS On-Orbit Status 27 Jan 2003

By SpaceRef Editor
January 27, 2003
Filed under , ,
ISS On-Orbit Status 27 Jan 2003

All ISS systems continue to function nominally, except as noted previously
or below. Week 9 for Expedition 6 is underway.

After station inspection and morning hygiene, before breakfast and first
exercise, all crewmembers underwent the regular periodic Russian biomedical
assessments MO-8 (body mass measurement, BMM) and MO-7 (calf volume measurement).
[FE-1 Nikolai Budarin set up the BMM mass measuring device, which uses
calibrated springs to determine the subject’s mass in weightless space,
and stowed it away after the tests. Calf measurements (left leg only)
are taken with the ISOG device, a custom-sewn fabric cuff that fits over
the calf, using the knee and lower foot as fixed reference points, to
provide a rough index of deconditioning in zero-G and effectiveness of

Later in the day, it was again time for Nikolai to conduct a session
of the MO-5 MedOps protocol of cardiovascular evaluation during graded
exercises on the VELO cycle ergometer, assisted by FE-2/SO Don Pettit
as CMO (crew medical officer). [Using the Gamma-1 ECG equipment with
biomed harness, skin electrodes, and a blood pressure and rheoplethysmograph
cuff wired to the cycle ergometer’s instrumentation panels, the cosmonaut
worked the pedals after a prescribed program at load settings of 125,
150, and 175 watts for three minutes each. All measurements were recorded
and telemetered to MCC-M, from where the workout was controlled by a

CDR Kenneth Bowersox continued his job of loading the new R.3 software
version on the EXPRESS racks,- today on ER3, the fourth of five ER upgrades.
[The loading of ER3, which is the first of two racks equipped with the
ARIS (active rack isolation system), involved the ELC (EXPRESS laptop
computer), RIC (rack interface controller) and ARIS displays, all from
a compact disk (for the previous racks Sox used the floppy drive). Test
of the ER3 ARIS with the new s/w will begin later this week, and to avoid
disturbances during the test, the crew today relocated some HRF (human
research facility) hardware items needed while the test is in process.]

FE-1 Nikolai Budarin continued the renal (kidney) stone prevention experiment
on his second day of the regimen, which required him to collect urine
samples in the course of the day besides diet logging. For Bowersox,
sample collection ended this morning, Pettit’s yesterday.

Don set up the hardware for EVARM-IV (EVA radiation monitoring no. 4),
to take on-board readings about 15 minutes after activating the badge
reader, then powered it down again. [EVARM badges will be used every
other week for the remainder of Increment 6 to characterize the IV-CPDS
(intravehicular charged particle directional spectrometer) environment,
and today’s procedure relocated the badges to the IV-CPDS area.]

Budarin conducted the periodic functional open/close test of the spare
emergency vacuum valve (AVK) of the COA (= ARS, atmosphere purification
system) Vozdukh carbon dioxide (CO2) removal system. He then restowed
the valve assembly behind a Service Module (SM) panel (321). [The AVKs
are critical because they close the Vozdukh’s vacuum access lines in
the event of a malfunction of the regular vacuum valves (BVK) or a depressurization
in the Vozdukh valve panel (BOA). Access to vacuum is required to vent
CO2 during the regeneration of the absorbent cartridges (PP). During
nominal operation, the AVK valves remain open.]

Budarin also continued the checkout of the recently removed Vozdukh
BVK-1 vacuum valve package. His preliminary test results showed it to
be OK, and more testing was scheduled today, including size verification
of a required special wrench.

The recent testing of the two removed Vozdukh VN vacuum pumps confirmed
that both are failed. They will be discarded on Progress 9P.

Later, Nikolai conducted a periodic comparison check of pressure gauge
readings, using four portable manometers (MV) from the SM to take readings
in the Soyuz spacecraft, to verify the Soyuz pressure gauges. [Acceptable
tolerance: plus/minus 2 mmHg discrepancy.]

Budarin also performed his regular daily inspection of the Russian BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

Bowersox completed the daily routine payload status checkup in the Lab
(PCG-STES010) and prepared the IMS inventory delta file for downlink,
while Pettit performed the routine maintenance of the SOSh life support
system in the SM.

At 8:40am, ISS attitude control was handed over the Russian MCS (motion
control system) thrusters for an attitude test of the onboard "Kurs" system
docking program by MCC-M. Later, the station was slewed to a biased docking
attitude, then briefly moded to Inertial (free drift) to support testing
of Progress vehicle docking with the hand-controlled TORU backup teleoperator
approach & docking system, with SM and FGB solar arrays temporarily
in fixed (non-tracking) position. Attitude control returns to USOS CMG
momentum management later this afternoon.

At 12:36pm EST, a ship-to-ship call took place between the ISS crew
and STS-107/Columbia, where the Red team of Rick Husband, Laurel Clark,
Ilan Ramon and Kalpana Chawla participated (Blue team being on "sleep
shift"). Video of the Shuttle crew was received on the ground via
Ku-band, and the two crews conferred via S-band audio (space-to-ground
2, S/G-2). Good job by all! [Steps were taken to avoid any audio bleed-over
into the Russian segment where Progress docking testing was underway.]

MCC-H conducted more troubleshooting for checking out the failed heaters
of the ETCS (external thermal control system) loop B NTA (nitrogen tank

All crewmembers performed their regular daily physical exercise program.

The Node starboard hatch window was properly installed by the crew over
the weekend and successfully leak-checked. [The two-pane porthole, which
took the place of the old ECOMM (early comm) antenna mounting plate,
now permits viewing into the Joint Airlock from the Node during EVAs.]

A three-page manifest of hardware and trash items to be loaded on Progress
M1-258 (9P) for disposal has been uplinked, along with detailed loading
instructions. 9P will undock on 2/1 (Saturday). [For reasons of flight
stability and control of the automated (pre-programmed) drone, loading
must be done with great care in choreographed stages according to a topological
plan that subdivides the Progress interior in planes and tiers. Proper
distribution of mass is mandatory to ensure acceptable center-of-gravity
location and mass moments of inertia. Thus, for example, containers are
stowed aboard in sequence, not in parallel. The loading process is documented
with a video survey, and the recording is inspected by specialists at
TsUP/MCC-M before Progress close-out is approved.]

Today’s targets for the CEO (crew earth observations program) were Perth,
Australia (looking a touch left for this coastal conurbation), Mekong
River Delta (large deltas are changing ever faster as agriculture upstream
expands in area. Agricultural disturbance of soil releases topsoil into
rivers, with consequences that are only just beginning to be understood.
We do know that delta islands change shape and grow as these processes
continue to operate. Detailed views of delta islands are thus requested),
Ho Chi Minh City, Vietnam (looking right of track on the easternmost
branch of the Mekong River delta about ten miles in from the sea), Cape
Town, South Africa (looking left for contextual views of this port city
at the north end of the Cape Peninsula), Industrialized SE Africa haze
(high atmospheric pressure is setting in so that haze should be visible
both sides of track. Southern Africa is the greatest generator of haze
in the southern hemisphere. Obliques to the right may reveal vertical
stratification, thicker haze over the plateau and thinner haze over the
Lesotho Highlands), Johannesburg, South Africa (looking mainly left of
track for views of the central zone of this 150-mile-long megacity),
Patagonian Glaciers (nadir views of glacier tongues were requested),
and High Central Andean Glaciers (looking left and right for any ice-capped
volcanoes. These tropical ice caps are melting at unprecedented rates
under present conditions.

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:45pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (16 Amp mode). Vozdukh CO2 scrubber
is On. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control
subsystem) is off. MCA (major constituents analyzer) is operating.
BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2
in Purify mode. RS air conditioner SKV-1 is Off; SKV-2 is Off.

SM Working Compartment: Pressure (mmHg) — 753; temperature (deg C)
— 24.3; ppO2 (mmHg) — 160.9; ppCO2 (mmHg) — 2.5 (suspect).
SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C)
— 19.3.
FGB Cabin: Pressure (mmHg) — 752; temperature (deg C) — 21.7.
Node: Pressure (mmHg) — 741.34; temperature (deg C) — 22.9 (shell);
ppO2 (mmHg) — 163.8; ppCO2 (mmHg) — 5.8.
U.S. Lab: Pressure (mmHg) — 743.91; temperature (deg C) — 22.6; ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 744.01; temperature (deg
C) — 21.8; shell heater temp (deg C) — 21.9, ppO2 (mmHg) — n/a; ppCO2
(mmHg) — n/a.
PMA-1: Shell heater temp (deg C) — 23.5
PMA-2: Shell heater temp (deg C) — 17.9
(n/a = data not available)

Propulsion System (PS): Total propellant load available (SM + FGB +
Progress) — 3669 kg (8089 lb) [as of 1/23/03]. (Capability: SM — 870
kg; FGB — 6160 kg).

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and
BGA 4B both in dual-angle mode (directed position).
SM batteries: Battery #7 is off line (failed); all other batteries (7)
are in "Full Charge" mode.
FGB batteries (3am): Battery #2 is offline; all other batteries (5) are
in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:
C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
EXT-1 is On (primary), EXT-2 is off.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is On (primary); PL-1 MDM is off (diagnostic
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane
1 is down (as of 11/14/02).
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line (CMG-1 failed).
State vector — U.S. SIGI-1
Attitude — U.S. SIGI-1
Angular rates — U.S. RGA-1

Flight Attitude:
LVLH TEA (local vertical/local horizontal = "earth-fixed":
z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch:
-9.1 deg, roll: 0 deg]), CMG/Thruster Assist Momentum Management).
Solar Beta Angle: 25.3 deg (magnitude decreasing).

Communications & Tracking Systems:
FGB MDM-1 is powered Off; FGB MDM-2 is operating.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 based at MBS PDGF2 with Keep Alive (KA) power on both
MBS: KA power on both strings.
MT: latched at WS4, with KA power.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 8:18am EST [= epoch]):
Mean altitude — 388.2 km
Apogee — 392.0 km
Perigee — 384.4 km
Period — 92.32 min.
Inclination (to Equator) — 51.63 deg
Eccentricity — 0.0005623
Orbits per 24-hr. day — 15.60
Altitude loss — 120 m (mean) in last 24 hours
Solar Beta Angle — 25.3 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 23906

For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.